OpenStudy (anonymous):

Evaluate tan 30° without using a calculator by using ratios in a reference triangle. What is the first thing to do here?

jimthompson5910 (jim_thompson5910):

do you have a 30-60-90 triangle with you?

OpenStudy (anonymous):

I know the 30-60-90 triangle.

jimthompson5910 (jim_thompson5910):

ok draw one out and tell me what is opposite the 30 degree angle

OpenStudy (anonymous):

wait it's 1

jimthompson5910 (jim_thompson5910):

it should be 1, unless you are working with the unit circle

jimthompson5910 (jim_thompson5910):

what is the adjacent side near the 30 degree angle

OpenStudy (anonymous):

yeah it's 1 since 90=2 and 60=sqr3

jimthompson5910 (jim_thompson5910):

the adjacent side should be sqrt(3)

jimthompson5910 (jim_thompson5910):

now divide opposite over adjacent

OpenStudy (anonymous):

it's 1/sqrt(3). :)

jimthompson5910 (jim_thompson5910):

Correct \[\Large \frac{1}{\sqrt{3}}\] or \[\Large \frac{\sqrt{3}}{3}\]

OpenStudy (anonymous):

:) Thanks!

jimthompson5910 (jim_thompson5910):

no problem